What is the probability that a yellow marble is drawn at random from a bag containing 3 white, 2 yellow, 1 green, and 4 blue marbles. (Example of how to enter answer: if the answer was two thirds you would enter 2/3)
Answer by htmentor(789)

What is the probability that a yellow marble is drawn at random from a bag containing 3 white, 2 yellow, 1 green, and 4 blue marbles. (Example of how to enter answer: if the answer was two thirds you would enter 2/3)
=======================
There are a total of 10 marbles in the bag, 2 of which are yellow
Therefore the P(yellow) = 2/10 = 1/5